OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4)

OpenStudy (anonymous):

its -5 f(4)= 2(4)-2=6 g(x)=3(4)-1=11 So... 6-11=-5